# Inferring flavor mixtures in multijet events

### v2

@@ 3.3.1 Results
To quantify and analyze the convergence to the true $c$- and $b$-distributions in GN1 for all models and also 1D and 4D scenarios, we have summarized the results in Fig. 7. For each run, we plot on the vertical axis the root mean square (RMS) distance between the true curve and the posterior samples averaged in all bins and classes. On the horizontal axis we compute the density on the true curve averaged for all bins and classes, plotting the exponential of the mean of the logarithm of this probability, see Ref. [9] for details. As it can be seen from Fig. 7, the posterior convergence improves as $N$ increases and as we use models that include more prior knowledge; in this case the continuity. This plot summarizes and quantifies the posterior convergence for the $c$- and $b$-distributions for all models and scenarios.
@@
Considering the posterior for the fractions of each one of the classes $cccc$, $ccbb$ and $bbbb$, we should read the right column in Figs. 36. We can see that in all cases the posterior distribution for the fractions is close to the true value in absolute distance, which is a good indicator. If one analyzes how it changes the probability in the true bin from the prior to the posterior, we also see that in all cases it increases. This quantifies the improvement in the posterior probability compared to the prior probability in the true fraction bin. We collect all these results in Fig. 8. Despite some statistical fluctuations, we also observe a pattern that corresponds to more probability in the true fraction as one increases the number of events.
@@
Figure 9: Summary results in distributions and fractions for three different random seeds. Plots are similar to Fig. 7 and 8, but for different data realizations because of changing the random seed. We observe the same general pattern.
Refer to captionRefer to caption
To bolster the presented results we show in Fig.9 the same summary plots for a few other runs using different seeds. We find that essentially the same pattern is observed. All the presented runs have specific seeds and can be reproduced with the information and scripts in [9].
In Fig. 10 we perform a light statistical analysis to compare the model behavior in retrieving the correct fraction for the classes in the dataset. We do not include the Point Estimate model since its behavior is very similar to Unimodal, from which it is derived. We study in the left panel of this figure the model robustness to avoid label-switching, and in the right panel its robustness to increase the posterior density in the true bin. We find that the Unimodal model may lay its posterior mean shifted from the true fraction (left panel), however its increase in probability at the true bin is as good as the other models (right panel). Dirichlet and Gaussian process models have a good centering of their posteriors around the true fraction and mostly without label-switching.
4 Discussion
The results in this article are connected to the results in Ref. [18], in which it is shown how Bayesian techniques generalize and improve some customary HEP tools for data analysis. In the present work, we have extended the aforementioned analysis by generalizing that the jets $b$-tagging score probability distribution can have arbitrary continuous, preferably unimodal, distributions. This result implies an important approach to the application of the proposed method in realistic analyses.
The current article, as Reference [18], are both building blocks that aim to construct a different approach to measure $pp\to hh\to b\bar{b}b\bar{b}$ by fully exploiting the available information in the data. Of course, this is a large enterprise that needs many building blocks. Further improvements are in the pipeline and include the treatment of some of the systematic uncertainties that may yield dependence among the in-principle conditionally independent observables. Progress in including systematic correlations in a mixture model can be found in Ref. [19], whose scheme could be transferable to this case. For a realistic analysis, all systematic uncertainties need to be studied, understood and modeled, and all relevant backgrounds included in the model. The most challenging background is non-resonant $b\bar{b}b\bar{b}$, because of its irreducibility in the jets flavor. Observe that despite that in current ATLAS [20] and CMS [21] $pp\to hh\to b\bar{b}b\bar{b}$ analyses this is the only main background, within the proposed Bayesian framework more backgrounds could play an important role in the analysis. This is because the proposed Bayesian techniques are such that many signal events that are tagged as $3b$ in the non-Bayesian framework, will enter into the analysis because of the soft-assignment enhancement (see discussion in Ref. [18]) that allows the inclusion of more signal events in the analysis.
Finally, and in general, one should envisage that the presented techniques that yield good inference in mixture models with arbitrary, continuous and preferably unimodal distributions, could play important enhancements in other areas. For instance, in observables where QCD corrections play an important role, one could start with the current order computation of some given distributions, and let the algorithm infer the true distributions for each class by seeing the data. Such a project on QCD corrections could be validated by starting with LO distributions and verifying whether the inference approaches the NLO or further order available distributions. Such a work would be favored if the process has independent observables that can provide multidimensionality to the problem.
5 Conclusions
We have implemented four models to infer binned $b$-tagging score distributions and the fraction of each class in a dataset consisting of $N=100$, $250$ and $500$ events with four jets per event. The allowed classes correspond to the $cccc$, $ccbb$ and $bbbb$ flavor combinations. In a Dirichlet model, we find that the lack of correlation in neighboring bins negatively affects the inference of the $c$- and $b$-distributions, whereas the inference for the fraction yields a result relatively close to the true fraction value. In a Gaussian process model, where continuity is exploited through the multivariate normal covariance matrix, we find that $c$- and $b$-distributions and the fraction are correctly extracted through the presented algorithm. We have developed a Unimodal model that samples a mixture of strictly unimodals weighted through a Dirichlet with small parameters, which yields mostly unimodal distributions. This model also correctly extracts the $c$- and $b$-distributions and the fraction. Using the Unimodal model as prior we have defined a Point-Estimate model which is strictly unimodal and also performs correct extraction of the relevant distributions and fraction in the problem. We summarize the performance of the models in Figs. 7 and 8, and we also show the summarized results for other seeds in Fig. 9. We find in all cases the same general pattern: inference works correctly as described above, and the posterior for the individual distributions and the fraction approaches the true value as the number of events $N$ increases. We stress that this is achieved by starting with a biased or agnostic prior and just seeing the data with the correct model for it, demonstrating that the method could be robust to extracting the correct distributions and fractions in a real scenario in which one also starts with a prior knowledge that in general does not match the true values. This result is useful to reduce the impact of Monte Carlo simulations, as well as to eventually reduce the need for calibrated $b$-tagging working points.
Acknowledgments
Appendix A Computational considerations

### v3

@@ 3.3.1 Results
To quantify and analyze the convergence to the true $c$- and $b$-distributions in GN1 for all models and also 1D and 4D scenarios, we have summarized in Appendix A the results in Fig. 10. For each run, we plot on the vertical axis the root mean square (RMS) distance between the true curve and the posterior samples averaged in all bins and classes. As a metric to assess a density estimation on the true values we proceed as follows. For each run, we sample datasets from the posterior with equal number of events as in the data, and from there we compute the density on the true curve averaged for all bins and classes, plotting the exponential of the mean of the logarithm of this probability. See Ref. [8] for details on these metrics. As it can be seen from Fig. 10, the posterior convergence improves as $N$ increases and as we use models that include more prior knowledge; in this case the continuity. This plot summarizes and quantifies the posterior convergence for the $c$- and $b$-distributions for all models and scenarios.
@@
Considering the posterior for the fractions of each one of the classes $cccc$, $ccbb$ and $bbbb$, we should read the right column in Figs. 36. We can see that in all cases the posterior distribution for the fractions is close to the true value in absolute distance, which is a good indicator. If one analyzes how it changes the probability in the true bin from the prior to the posterior, we also see that in all cases it increases. This quantifies the improvement in the posterior probability compared to the prior probability in the true fraction bin. We collect all these results in Appendix A in Fig. 11. Despite some statistical fluctuations, we also observe a pattern that corresponds to more probability in the true fraction as one increases the number of events.
Refer to captionRefer to caption
To bolster the presented results we show in Appendix A the Fig.12 with the same summary plots for a few other runs using different seeds. We find that essentially the same pattern is observed. All the presented runs have specific seeds and can be reproduced with the information and scripts in [8].
In Fig. 7 we perform a exploratory analysis to compare the model behavior in retrieving the correct fraction for the classes in the dataset. We do not include the Point Estimate model since its behavior is very similar to Unimodal, from which it is derived. We study in the left panel of this figure the model robustness to avoid label-switching, and in the right panel its robustness to increase the posterior density in the true bin. We see in the left panel that the Unimodal model may have its posterior mean (cyan points) shifted from the true fraction (magenta cross), however we see in the right panel that its increase in probability at the true bin is as good as the other models. Dirichlet and Gaussian process models have a good centering of their posteriors around the true fraction and mostly without label-switching.
Finally, in order to test the inference power in more realistic situations in which the signal class—usually $bbbb$—has a smaller fraction, we also run the inference model in samples with 10% and 1% $bbbb$ fractions. We present these results in Fig. 8 for $N=500$ and using the Unimodal model (the full plots dataset as in Figs. 36 can be found in Ref.[8]). We observe that also in these cases the inference runs correctly in the distributions (not shown) and more important in the small fractions. We find that the background class $ccbb$, whose fraction is not small, is playing a key role to provide information on the individual $b$-component distribution, which is simultaneously used to infer the very small $bbbb$ fraction. It is worth also noting that we start with an agnostic uniform prior on the fractions, whereas in a real case one still has room to improve by using some prior information on the expected fractions for the classes.
3.4 Comparison to cut-based methods
In order to assess the potential convenience and advantages of the presented Bayesian method in real physics problems, it is useful to compare it to usual cut-based methods used at the LHC. To make a fair comparison, we compare the classification performance of (the usually) signal $bbbb$ events in our Bayesian and the cut method. We use for this comparison the sample with $N=500$ events and 10% $bbbb$ signal fraction. We consider this sample to have large backgrounds while still having enough signal events ($\sim 50$) to have smooth plots.
In the usual cut-based method, in order to have a high $bbbb$ purity sample, analyses select events by requiring that all four jets have a $b$-score above some threshold. This is the case, for instance, in the $pp\to hh\to b\bar{b}b\bar{b}$ analyses in Refs. [19] and [20]. Therefore, one can assess the performance of such a classifier by constructing a Receiving Operating Characteristic (ROC) curve as the mentioned threshold is varied. In each selected sample one has a different True and False Positive Ratios, which draws the so-called ROC curve.
In the Bayesian framework one has that events are modeled as being sampled from a distribution. In this case, being a mixture model, there is a latent variable $z_{n}$ whose outcome determines the class for the $n^{th}$ event (see for instance Refs. [21, 22]). One can easily compute the probability for this variable conditioned on the outcome of the whole data $X$, and in particular on its corresponding measured data $x_{n}$. Using Bayes theorem
$p(z_{n}|X)=p(z_{n}|x_{n},X_{/n})=\frac{p(x_{n}|z_{n},X_{/n})\,p(z_{n}|X_{/n})}{p(x_{n}|X_{/n})},$ (8)
where $X_{/n}$ includes all data but the $n^{th}$ event, although in the practice we approximate it by the whole data without causing any noticeable bias. The first factor in the numerator is easily computed using the posterior samples from the inference problem to numerically evaluate the following expression
[…]
The second factor in the numerator of Eq. 8 is obtained analogously, but integrating out on the parameters that estimate the fractions, whereas the denominator is the result of integrating out $z_{n}$ in the numerator. In this way, one can compute $\Pr(z_{n}=bbbb|X)$ for each event and use this probability as a parameter to define a Bayesian tagging score for the $bbbb$ class. We can use a threshold on $\Pr(z_{n}=bbbb)$ to select events in the sample and therefore a ROC curve that evaluates this classifier.
The ROC curves indicates the classification ability to identify the $bbbb$ events. We show these curves from our Bayesian and the cut method as the solid-blue and solid-red lines, respectively, in the left panel in Fig. 9, where it can be seen that the Bayesian classifier outperforms the cut-based one. Moreover, the Bayesian framework can also work as a classifier for the other classes in the problem ($ccbb$ and $cccc$). We show the classifying performance for those classes as well with their corresponding ROC curves in the right panel in Fig. 9.
Observe that the aforementioned ROC curves require the true labels of the events. However in a realistic situation, we do not have true labels and we can only estimate ROC curves. It is interesting to study the level of agreement between the estimated ROC curve and the true ROC curve. In the cut-based method the estimated curve is computed from the priors of the individual $c$- and $b$-components by calculating the probability of having four $b$-tagged for a given threshold in the tagging score. Using the priors mean for the $c$- and $b$-components from the previous section, this yields the dashed-red curve in the left panel in Fig. 9. Whereas for the Bayesian framework one can do as follows: since one has a posterior on all the parameters, one can simulate new synthetic data with labels and then compute a Bayesian ROC curve using labeled synthetic data [^4] Another naive possibility is to use the probability of each event to belong to each class as a soft-assignment to estimate the number of events from each class in each selection. However, this method estimates number of events using the same variable that one varies to select events for the ROC curve, and therefore inducing an expected bias in favor of an overestimated ROC curve. We thank M.Szewc for this observation.. This ROC curve is plotted as dashed-blue in the same plot. Observe, therefore, that the ROC curve estimation has better agreement in the Bayesian framework than in the cut-based one. This is also an important result, since a biased ROC curve induces biased results. Also notice that the Unimodal inference model only uses as prior information unimodality, continuity and that in the first bins the individual $c$-component has larger probability than the $b$-component, to avoid label switching. Thus in case that the priors mean on the $c$- and $b$-components varies, the blue lines remain the same, wheres the dashed-red will shift along the ROC plane. We should remark that the disagreement between dashed-red and solid-red originates in the discrepancies between prior and true distributions for the $c$ and $b$ individual components, which is one of the central problems addressed with the methods presented in this work.
Refer to captionRefer to caption
Figure 8: Case with 10% (left) and 1% (right) $bbbb$ signal fraction in the sample to showcase the inference power still with a small signal fraction. The results are inferred through the Unimodal model and correspond to $N=500$ events and an agnostic flat prior on the fractions, as in the previous figures. Interestingly, the inference of a small fraction in the $bbbb$ signal class is assisted through the $ccbb$ background class which, instead of being an issue, becomes a leverage to infer the individual $b$-component and therefore, simultaneously, the $bbbb$ fraction. The complete set of results for these fractions can be found in Ref. [8].
Refer to captionRefer to caption
Figure 9: Left: Comparison of ROC curves between the usual cut-based method (red) and the presented Bayesian method (blue) as classifiers for $bbbb$ events. In the usual cut-based method one selects $bbbb$ events by requiring that all jets have a $b$-score larger than a given threshold, whereas in the Bayesian method one has a posterior for a latent variable for each event, $z_{n}$, which indicates the probability of each event to belong to each class. Solid curves represent the true ROC curves using the true labels for each event, whereas dashed lines represent the more realistic case in which one does not have the true label for each event (see text). The solid-blue being above the solid-red indicates that the Bayesian method is a better classifier. The agreement between the dashed-blue and solid-blue (in comparison to the red ones) indicate that the Bayesian method is less biased in estimating the ROC curves. Right: As a byproduct of the Bayesian method, one can also classify the events in each one of the available classes using $z_{n}$. Here we show the classification power of the method for each one of the classes through their respective ROC curve. (In both plots we use the Unimodal model with $N=500$ events and 10% $bbbb$ fraction)
4 Discussion
With respect to the 4D examples, it is worth analyzing what is leveraging the inference results. There are a few features that should be mentioned. On one hand, being the events four-dimensional and the four scores conditionally independent, reduces considerably the number of unknowns. Here conditionally independence means that once the class has been assigned ($cccc$, $ccbb$ or $bbbb$), each $b$-tagging score is sampled independently of the outcome of the others; there is no covariance matrix to be inferred. On the other hand, the four conditionally independent observables in each one of the three classes, come only from two distributions, the $c$- and $b$-individual components. This reduces still more the number of unknowns.
The results in this article are connected to the results in Ref. [22], in which it is shown how Bayesian techniques generalize and improve some customary High Energy Physics tools for data analysis. In the present work, we have extended the aforementioned analysis by generalizing that the jets $b$-tagging score probability distribution can have arbitrary continuous, preferably unimodal, distributions. This result implies an important approach to the application of the proposed method in realistic analyses.
The current article, as Reference [22], are both building blocks that aim to construct a different approach to measure $pp\to hh\to b\bar{b}b\bar{b}$ by fully exploiting the available information in the data. Of course, this is a large enterprise that needs many building blocks. Further improvements are in the pipeline and include the treatment of some of the systematic uncertainties that may yield dependence among the in-principle conditionally independent observables. Progress in including systematic correlations in a mixture model can be found in Ref. [23], whose scheme could be transferable to this case. For a realistic analysis, all systematic uncertainties need to be studied, understood and modeled, and all relevant backgrounds included in the model. The most challenging background is non-resonant $b\bar{b}b\bar{b}$, because of its irreducibility in the jets flavor. Observe that despite that in current ATLAS [24] and CMS [25] $pp\to hh\to b\bar{b}b\bar{b}$ analyses this is the only main background, within the proposed Bayesian framework more backgrounds could play an important role in the analysis. This is because the proposed Bayesian techniques are such that many signal events that are tagged as $3b$ in the non-Bayesian framework, will enter into the analysis because of the soft-assignment enhancement (see discussion in Ref. [22]) that allows the inclusion of more signal events in the analysis. Current analyses for the proposed $pp\to hh\to b\bar{b}b\bar{b}$ channel using fixed $b$-tag working points, with efficiency 0.77 as in [24], end up loosing a $1-0.77^{4}$ part of the signal events (i.e. 65% of the $bbbb$ are discarded). Another point that should be considered in expanding the present analysis for a realistic application is that many times a jet can lie beyond the tagging limit (usually $|\eta|=2.5$), which in the practice would break the hypothesis that flavors come in pairs. Although this is unlikely for one jet, having four jets makes it more likely. A possibility to overcome this difficulty without neglecting these events, consists in using also in the model the invariant masses of the jet pairs, as indicated in [22]. In this way we can assign in the likelihood a probability distribution for a jet that lies beyond $|\eta|=2.5$ to be flavor-paired with another jet by studying their invariant mass within the knowledge of the system under study. As a final observation of improvements that should follow the present results we should mention that in a real study one should perform a posterior predictive check (see for instance [26, 21]). This study computes the probability of the data to have been sampled from the proposed model using the inferred parameters. Therefore, being a measure of the compatibility between the model and the data.
Finally, and in general, one should envisage that the presented techniques that yield good inference in mixture models with arbitrary, continuous and preferably unimodal distributions, could play important enhancements in other areas. For instance, in observables where QCD corrections play an important role, one could start with the current order computation of some given distributions, and let the algorithm infer the true distributions for each class by seeing the data. Such a project on QCD corrections could be validated by starting with LO distributions and verifying whether the inference approaches the NLO or further order available distributions. Such a work would be favored if the process has independent observables that can provide multidimensionality and conditionally independence to the problem.
5 Conclusions
We have implemented four models to infer binned $b$-tagging score distributions and the fraction of each class in a dataset consisting of $N=100$, $250$ and $500$ events with four jets per event. The allowed classes correspond to the $cccc$, $ccbb$ and $bbbb$ flavor combinations. In a Dirichlet model, we find that the lack of correlation in neighboring bins negatively affects the inference of the $c$- and $b$-distributions, whereas the inference for the fraction yields a result relatively close to the true fraction value. In a Gaussian process model, where continuity is exploited through the multivariate normal covariance matrix, we find that $c$- and $b$-distributions and the fraction are correctly extracted through the presented algorithm. We have developed a Unimodal model that samples a mixture of strictly unimodals weighted through a Dirichlet with small parameters, which yields mostly unimodal distributions. This model also correctly extracts the $c$- and $b$-distributions and the fraction. Using the Unimodal model as prior we have defined a Point-Estimate model which is strictly unimodal and also performs correct extraction of the relevant distributions and fraction in the problem. We summarize the performance of the models in Figs. 10 and 11, and we also show the summarized results for other seeds in Fig. 12. We find in all cases the same general pattern: inference works correctly as described above, and the posterior for the individual distributions and the fraction approaches the true value as the number of events $N$ increases. We stress that this is achieved by starting with a biased or agnostic prior and just seeing the data with the correct model for it, demonstrating that the method could be robust to extracting the correct distributions and fractions in a real scenario in which one also starts with a prior knowledge that in general does not match the true values. This result is useful to reduce the impact of Monte Carlo simulations, as well as to eventually reduce the need for calibrated $b$-tagging working points.
Acknowledgments
Appendix A Additional results as plots
Additional plots (Figs. 10-12) to better visualize and understand the scope of the proposed method.
@@ Appendix A Additional results as plots
Figure 12: Summary results in distributions and fractions for three different random seeds. Plots are similar to Fig. 10 and 11, but for different data realizations because of changing the random seed. We observe the same general pattern.
Appendix B Computational considerations